# A Method for On‑Orbit Calibration of the VLAST‑P Electromagnetic Calorimeter

### v1

@@ 4.1 Simulation Methodology
Cosmic ray events consist of about 90% protons, 10% helium nuclei, and other heavy ions; unlike the sea-level cosmic rays, the main component of which is the muon. Thus, calibration of the sensitive units in space-based detectors is most commonly performed using minimum ionizing protons from spatial cosmic rays. This simulation aims to investigate the MIP-induced signal response in a 5 × 5 crystal array detector with GEANT4, through the development of a backtracing database grounded in a realistic geomagnetic field model.
Backtracing is generally easier and more efficient than forward tracing, as it launches from the observation point and avoids sampling the vast incident phase space. A comprehensive global geomagnetic backtracing database has been constructed using specialized software that simulates particle trajectories through an eighth-order Runge–Kutta numerical integration method within the International Geomagnetic Reference Field (IGRF-13) model named the GeoMagFilter database [11]. The database is constructed with three parameters: particle rigidity, zenith angle, and azimuth angle. Particle rigidity is defined as momentum divided by the charge quantity of the particle to represent how difficult it is for a particle to be deflected by a magnetic field. The latter two parameters can be calculated based on geographic longitude and latitude, which are typically accessible as basic orbital information of the satellite. The sampling grid spans longitudes from $0^{\circ}$ to $360^{\circ}$ in $10^{\circ}$ intervals and latitudes from $75^{\circ}$ to $75^{\circ}$ in $5^{\circ}$ steps, covering all possible orbital positions of both the International Space Station (ISS) and the China Space Station (CSS). Directional angles are sampled uniformly on a spiral in $4\pi$ steradians with an angular spacing of 0.03 radians, resulting in approximately 13,963 distinct directions. A binary value is assigned to indicate whether a particle can successfully trace back beyond the geomagnetic field (1 for successful backtracing, 0 for failure, e.g., when the trajectory terminates within $10^{5}$ km).
The assumed orbital altitude is 500 km, being consistent with the orbital altitude of the VLAST-P satellite, and the database will tell which rigidity a particle can be backtraced successfully when the satellite lies at a certain position. By dividing the $4\pi$ steradians, distribution of proton flux of each distinct direction can be obtained. With a built-up dedicated all detectors simulation, one can randomly do sampling of particle incident direction and particle rigidity considered as cosmic ray source. In order to generate events efficiently, the following method is adopted in the simulation.
@@ 4.2 Event Selection
Due to arbitrary incident direction of cosmic ray particles, it is really necessary to distinguish the events that can be used for calibration. First, according to the trigger logic, during the MIP calibration process, it is required that the hit signals from multiple detectors be combined using a logical AND operation. This ensures that only events where several detectors are simultaneously triggered are selected for calibration. A more specific set of thresholds is applied in the simulation, requiring a logical AND signal between the two ACD layers, as well as valid energy depositions in both the converter region of the tracker and the calorimeter. An event is kept if at least two units in two ACD layers, one for each, have more than 0.6 MeV of deposited energy, one converter unit has more than 2.52 MeV, and one ECAL cell has more than 36 MeV. These thresholds help make sure that the signals come from real particle interactions rather than random background. The input spectrum used to simulate the realistic spatial cosmic ray environment is generated from the database described earlier. From these input spectra, a pronounced east–west effect is observed, as shown in figure 4.
@@
For some events, the particle penetrates the detector at a large incident angle relative to the transverse plane. It will generate plenty of hit signals in ECAL. Thus, after excluding background or noise events, a hit cell number cut has been implemented. The hit number is limited to less than 6, which makes the path length in each cell reasonable. Furthermore, a correction for the path length has been applied in this step by evaluating the true trajectory using the incident angle and initial position of the particle. There lies a tracker in front of the ECAL. In the real on-orbit calibration, with the reconstruction algorithm of the tracker, the incident angle can also be used to select events. Additionally, in some events, it is inevitable that some hits come from a very short path length. To exclude these, hits with a path length less than 2 cm are not counted.
@@ 4.3 Performance
The selection criteria are primarily designed to reject events caused by particles entering at large incident angles, thereby preserving genuine events for calibration purposes. As presented in table 1, each selection stage leads to a substantial reduction in the number of surviving events. The table also lists the selection efficiency at each stage and the cumulative efficiency up to that point. Starting from the full simulated dataset, the event count decreases step-by-step, ultimately retaining only 2.95% of the total events with a source filter mentioned in Section 4.1. This result demonstrates that the selection strategy is effective in suppressing background while preserving a clean sample of physical events suitable for further analysis.
@@
On orbit, the satellite can operate in a sun-pointing mode, in which the front side of the detector continuously faces the Sun. Under this configuration, the detector alternately points toward the west and east directions with an approximately equal probability. Since the incident particle rigidity spectrum can differ significantly between the two orientations, we investigate this effect. The analysis indicates that the resulting rigidity distribution exhibits a discrepancy of less than 6%, as illustrated in figure 5.
@@
In the actual space environment, secondary particles, most of which are protons, also exist under space conditions, typically with energies below $10\,\mathrm{GeV}$. These secondary particles originate from interactions between high-energy primary cosmic rays and atmospheric atoms, producing cascades of new particles through nuclear reactions. According to the AMS proton flux measurements, the integral flux of secondary protons accounts for approximately 40% of the total integral flux of all protons in the equatorial region [12]. This component has also been accounted for in the reported results. The secondary events are combined with primary protons in proportion to their respective integral fluxes. The on-orbit MIP calibration results show that the energy deposition spectrum of protons follows a Landau-like distribution, with the most probable value (MPV) measured at $112.9\,\mathrm{MeV}$ in a single CsI(Tl) unit, as shown in figure 6(a). In the case of helium nuclei (He), it should be emphasized that their flux in space is also considerable, making them suitable for detector calibration. Compared with protons, helium nuclei are almost free from secondary components, as the production of secondary helium through nuclear interactions is extremely rare. A comparable distribution is obtained for helium nuclei, whose MPV is $443.4\,\mathrm{MeV}$, nearly four times that of the proton MIP, depicted in figure 6(b). This scaling behavior can be naturally understood from the Bethe–Bloch formalism, where the stopping power of charged particles in matter is approximately proportional to the square of their charge number ($Z^{2}$). As helium nuclei carry twice the charge of protons ($Z=2$), their energy deposition is therefore expected to be enhanced by a factor of about four, in good agreement with the observed ratio. Furthermore, observations suggest that secondary cosmic rays mainly come from regions near the sources and from the general interstellar medium [13]. This indicates that other products of cosmic-ray interactions in these regions, such as antiprotons, positrons, and spallation-produced Li, Be, and B, can also be detected in cosmic rays. Here, we focus on the flux of secondary protons in the equatorial region.
@@ 5 Conclusion And Outlook
To perform calibration in space, we developed a database based on geomagnetic backtracing to understand what properties charged particles must have to reach the satellite from outer space. Using this information, we created a realistic environment for simulating minimum ionizing protons, which are used to calibrate each ECAL channel. We also applied a set of selection criteria to ensure that the events used for calibration are clean and reliable. These criteria include trigger logic selection, hit number cuts, and path length corrections. The results show that, following the application of our calibration procedure, the proton and helium MIP spectra can be accurately calibrated with a spherical incident particle source, with an efficiency of approximately 2.95%. By additionally applying the path length correction, the MIP spectra can be accurately adjusted to their original appearances. This approach ultimately produces well-calibrated MIP spectra for all channels, confirming the effectiveness of the method.
To quantify the on-orbit calibration time requirement, the incident proton flux was taken as $\Phi_{p}=0.015\ \mathrm{cm^{-2}\ s^{-1}\ sr^{-1}}$, representative of the near-equatorial geomagnetic cutoff conditions. For each orbital revolution, an effective observation window of approximately $t_{\mathrm{eff}}=1200\ \mathrm{s}$ is available within the target geomagnetic latitude range ($20^{\circ}<\mathrm{lat}<20^{\circ}$) [14], based on an orbital period of $T_{\mathrm{orbit}}=1.5\ \mathrm{h}=5400\ \mathrm{s}$. The electromagnetic calorimeter has a geometric acceptance of $\Omega=6.34\times 10^{-3}\penalty 10000\ \mathrm{sr}$ for MIP calibration, corresponding to an event selection efficiency of $\varepsilon_{\mathrm{sel}}=0.0643\%$, as determined from Monte Carlo simulations employing a spherical source with a radius of 0.5 m. Consequently, 1142 events pass all calibration selection criteria in one orbit. A statistically robust calibration typically requires about $3\times 10^{3}$ events per channel across the 25 readout channels. In total, approximately $7.5\times 10^{4}$ events are required. Consequently, the total acquisition requirement amounts to approximately 65.7 orbital periods, corresponding to about 98 hours ($\approx 4.1$ days) of effective data-taking. This estimate serves as a practical baseline for scheduling and planning the calibration campaigns during the operational phase of the mission.
In conclusion, our calibration method is practical and effective. It ensures that VLAST-P can provide accurate and stable MIP calibration on orbit, even with randomly incident particle sources. MIP calibration plays a vital role in the energy measurement of gamma rays. Furthermore, this work lays the foundation for future improvements and applications in the full-scale VLAST mission.

### v2

@@ 4.1 Simulation Methodology
Primary cosmic rays consist of about 90% protons, 10% helium nuclei, and other heavy ions.This is in contrast to the cosmic ray flux at sea level, which is dominated by secondary muons Thus, calibration of the sensitive units in space-based detectors is most commonly performed using minimum ionizing protons from primary cosmic rays. This simulation aims to investigate the MIP-induced signal response in a 5 × 5 crystal array detector with GEANT4, through the development of a backtracing database grounded in a realistic geomagnetic field model.
Backtracing is generally easier and more efficient than forward tracing, as it launches from the observation point and avoids sampling the vast incident phase space [osti_4784046]. A comprehensive global geomagnetic backtracing database, named GeoMagFilter, has been constructed in Ref. [CHEN20255450]. It was generated using specialized software that simulates particle trajectories via an eighth-order Runge-Kutta numerical integration method within the International Geomagnetic Reference Field (IGRF-13) model. The database is constructed with three parameters: particle rigidity, zenith angle, and azimuth angle. Particle rigidity is defined as momentum divided by the charge quantity of the particle to represent how difficult it is for a particle to be deflected by a magnetic field. The latter two parameters can be calculated based on geographic longitude and latitude, which are typically accessible as basic orbital information of the satellite. The sampling grid spans longitudes from $0^{\circ}$ to $360^{\circ}$ in $10^{\circ}$ intervals and latitudes from $75^{\circ}$ to $75^{\circ}$ in $5^{\circ}$ steps, covering all possible orbital positions of both the International Space Station (ISS) and the China Space Station (CSS). Directional angles are sampled uniformly on a spiral in $4\pi$ steradians with an angular spacing of 0.03 radians, resulting in approximately 13,963 distinct directions. A binary value is assigned to indicate whether a particle can successfully trace back beyond the geomagnetic field (1 for successful backtracing, 0 for failure, e.g., when the trajectory terminates within $10^{5}$ km).
Given the orbital altitude of 500 km consistent with satellite, the database indicates the rigidity a particle can be backtraced successfully when the satellite is at a certain position. By dividing the $4\pi$ steradians, distribution of proton flux of each distinct direction can be obtained. Using a dedicated full-detector simulation, the incident direction and rigidity of particles are sampled to mimic a cosmic ray source. In order to generate events efficiently, the following method is adopted in the simulation.
@@ 4.2 Event Selection
Due to arbitrary incident direction of cosmic ray particles, it is necessary to distinguish the events that can be used for calibration. First, according to the trigger logic, during the MIP calibration process, it is required that the hit signals from multiple detectors be combined using a logical AND operation. This ensures that only events where several detectors are simultaneously triggered are selected for calibration. A more specific set of thresholds is applied in the simulation, requiring a logical AND signal between the two ACD layers, as well as valid energy depositions in both the converter region of the tracker and the calorimeter. An event is kept if at least two units in two ACD layers, one for each, have more than 0.6 MeV of deposited energy, one converter unit has more than 2.52 MeV, and one ECAL cell has more than 36 MeV. These thresholds help to ensure that the signals come from real particle interactions rather than random background. The input spectrum used to simulate the realistic spatial cosmic ray environment is generated from the database described earlier. From these input spectra, a pronounced east–west effect is observed, as shown in figure 4.
@@
For particles penetrating the detector at large incident angles, the path length calculation can be complicated by shower development. To determine the optimal selection criterion, we analyzed the hit cell multiplicity distributions for both showering and non-showering events. Based on this study, a threshold of fewer than 6 hit cells was established. This cut effectively discriminates against showering events, ensuring that the selected tracks have well-defined path lengths within the crystals. Furthermore, a correction for the path length has been applied in this step by evaluating the true trajectory using the incident angle and initial position of the particle. There lies a tracker in front of the ECAL. In the real on-orbit calibration, with the reconstruction algorithm of the tracker, the incident angle can also be used to select events. Additionally, in some events, it is inevitable that some hits come from a very short path length. To exclude these, hits with a path length less than 2 cm are not counted.
@@ 4.3 Performance
The selection criteria are primarily designed to reject events caused by particles entering at large incident angles, thereby retaining events suitable for calibration purposes. As presented in table 1, each selection stage leads to a substantial reduction in the number of surviving events. The table also lists the selection efficiency at each stage and the cumulative efficiency up to that point. Starting from the full simulated dataset, the event count decreases step-by-step, ultimately retaining only 2.95% of the total events with a source filter mentioned in Section 4.1. This result demonstrates that the selection strategy is effective in suppressing showering events while preserving a clean sample of physical events suitable for further analysis.
@@
On orbit, the satellite can operate in a sun-pointing mode, in which the front side of the detector continuously faces the Sun. Under this configuration, the detector alternately points toward the west and east directions with an approximately equal probability. Since the incident particle rigidity spectrum can differ significantly between the two orientations, the rigidity spectra shown in figure 4 are separately adopted as input distributions in the simulation to evaluate their impact. The analysis indicates that the MPV of the energy deposition for protons arriving from the East and West directions exhibits a discrepancy of approximately 6%, as illustrated in figure 5.
@@
In the actual space environment, secondary particles, most of which are protons, also exist under space conditions, typically with energies below $10\,\mathrm{GeV}$. These secondary particles originate from interactions between high-energy primary cosmic rays and atmospheric atoms, producing cascades of new particles through nuclear reactions. According to the AMS proton flux measurements, the integral flux of secondary protons accounts for approximately 40% of the total integral flux of all protons in the equatorial region [SecondaryProton]. In this study, secondary protons are modeled using the kinetic energy distribution measured by AMS, with arrival directions sampled isotropically from a source sphere. The simulation results show that the hit multiplicity distribution follows an approximately exponential decay, indicating that most secondary protons deposit energy in only a few crystals. The energy distribution of secondary protons is broader compared to that of primary protons. The selection efficiency of secondary protons is about half of the primary efficiency. These secondary events are combined with the primary proton events in proportion to their respective integral fluxes.
The on-orbit MIP calibration results show that the energy deposition spectrum of protons follows a Landau-like distribution, with the most probable value (MPV) measured at $112.9\,\mathrm{MeV}$ in a single CsI(Tl) unit, as shown in figure 6(a). In the case of helium nuclei (He), it should be emphasized that their flux in space is also considerable, making them suitable for detector calibration. Compared with protons, helium nuclei are almost free from secondary components, as the production of secondary helium through nuclear interactions is extremely rare. A comparable distribution is obtained for helium nuclei, whose MPV is $443.4\,\mathrm{MeV}$, nearly four times that of the proton MIP, depicted in figure 6(b). This scaling behavior can be naturally understood from the Bethe–Bloch formalism, where the stopping power of charged particles in matter is approximately proportional to the square of their charge number ($Z^{2}$). As helium nuclei carry twice the charge of protons ($Z=2$), their energy deposition is therefore expected to be enhanced by a factor of about four, in good agreement with the observed ratio. Furthermore, observations suggest that secondary cosmic rays mainly come from regions near the sources and from the general interstellar medium [SecondaryGeneral].
@@ 5 Conclusion And Outlook
To perform on-orbit calibration, we used a database based on geomagnetic backtracing to determine the cutoff rigidity required for charged particles to reach the satellite. Using this information, we created a realistic environment for simulating minimum ionizing protons, which are used to calibrate each ECAL channel. We also applied a set of selection criteria to ensure that the events used for calibration are clean and reliable. These criteria include trigger logic selection, hit number cuts, and path length corrections. The results show that, following the application of our calibration procedure, the proton and helium MIP spectra can be accurately calibrated with a spherical incident particle source, with an efficiency of approximately 2.95%. By additionally applying the path length correction, the MIP spectra can be accurately adjusted to their original appearances. This approach ultimately produces well-calibrated MIP spectra for all channels, confirming the effectiveness of the method.
To quantify the on-orbit calibration time requirement, the incident proton flux was taken as $\Phi_{p}=0.015\ \mathrm{cm^{-2}\ s^{-1}\ sr^{-1}}$, representative of the near-equatorial geomagnetic cutoff conditions. For each orbital revolution, an effective observation window of approximately $t_{\mathrm{eff}}=1200\ \mathrm{s}$ is available within the target geomagnetic latitude range ($20^{\circ}<\mathrm{lat}<20^{\circ}$) [DAMPE20], based on an orbital period of $T_{\mathrm{orbit}}=1.5\ \mathrm{h}=5400\ \mathrm{s}$. Based on Monte Carlo simulations employing a spherical source with a radius of $0.5\,\mathrm{m}$, the geometric factor of the electromagnetic calorimeter for MIP calibration is calculated to be $G=6.34\times 10^{-3}\penalty 10000\ \mathrm{m^{2}\cdot sr}$. Taking into account the proton flux and the orbital period mentioned above, this results in an expected yield of approximately 1142 valid calibration events per orbit. Considering the combined flux of primary and secondary protons, a sample of approximately $3\times 10^{3}$ events per channel is required across the 25 readout channels to constrain the statistical uncertainty to $0.5\%$. In total, approximately $7.5\times 10^{4}$ events are required. Consequently, the total acquisition requirement amounts to approximately 65.7 orbital periods, corresponding to about 98 hours ($\approx 4.1$ days) of effective data-taking. This estimate serves as a practical baseline for scheduling and planning the calibration campaigns during the operational phase of the mission.
In conclusion, we present an effective on-orbit calibration strategy. A key feature of our work is using a geomagnetic backtracing database instead of a standard random spherical source simulation. This makes the simulation more realistic by including accurate geomagnetic cutoffs for the incident particles. MIP calibration is critical for the accurate energy reconstruction of gamma rays. Moreover, this study establishes a solid framework for future applications in the full-scale VLAST mission.